OpenStudy (anonymous):

Find the sum of the first 8 terms of the sequence. Show all work for full credit. 1, -3, -7, -11, ...

OpenStudy (kmeis002):

Try to find a pattern connecting each number, start simple. Is there any number I can add/subtract from 1 to get -3? Does this same rule hold for -3 to -7?

OpenStudy (anonymous):

i see differences of -2, -4, -6. i see the pattern. but isn't there something else I have to do?

OpenStudy (anonymous):

a=1 d=-3-1=-4 It is an A.P. \[Sn=\frac{ n }{2 }\left\{ 2 a+\left( n-1 \right)d \right\}\] n=8, Put the values and get the result.

OpenStudy (kmeis002):

Once you have the pattern, you can come up with a rule. For this problem, if you subtract 3 from each term, you will get the next term: \[ (1-4) = -3, (-3-4) = -7, etc \] So our rule can be: \[ -3n+1 \] where n starts at zero, so, plug in n = 0 to 7 to get the first 8 terms

OpenStudy (kmeis002):

-4n+1 is the rule, sorry for the typo

OpenStudy (anonymous):

ohhh. i believe I understand it now. thanks :)